OpenStudy (anonymous):

write an equation for the line containing (-5,-16) and (-3,-8)

OpenStudy (anonymous):

\[m=\frac{ y^2-y^1 }{ x^2-x^1}\] \[m=\frac{ -8-(-16) }{ -3-(-5) }\]

OpenStudy (anonymous):

m=4

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

y=mx+b y=4x+b -8=4(-3)+b -8=-12+b b=4 y=4x+4

OpenStudy (anonymous):

ur welcome